OpenStudy (anonymous):

How many solutions does the equation −3a + 3a + 6 = 7 have? One Two Infinitely many None

OpenStudy (anonymous):

what school u go to

OpenStudy (anonymous):

flvs

OpenStudy (anonymous):

it has two i think i might have done the math wrong

OpenStudy (diamondboy):

actually its 1

OpenStudy (welshfella):

-3a + 3a = 3a - 3a = ?

OpenStudy (diamondboy):

wait...I am wrong @welshfella is right

OpenStudy (welshfella):

some value - itself = ?

OpenStudy (anonymous):

@connections8 look at her smart scor

OpenStudy (welshfella):

the 3a's disappear

OpenStudy (anonymous):

score

OpenStudy (anonymous):

yes cause -3 +3 =0

OpenStudy (anonymous):

what @lsherron doo u go ton concneciton 8

OpenStudy (welshfella):

so you are left with 6 = 7 there are no solutions

OpenStudy (anonymous):

okay thank you

OpenStudy (welshfella):

yw